OpenStudy (anonymous):

How many solutions within a unit circle \(|z| < 1\) does the equation \((1 + z)^{n + m} = z^n\) have for \(z\) complex, and \(n\), \(m\) positive integers?

ganeshie8 (ganeshie8):

@BSwan

ganeshie8 (ganeshie8):

@Kainui

ganeshie8 (ganeshie8):

\(\large \left(1+z\right)^m = \left(\dfrac{z}{1+z}\right)^n\)

OpenStudy (anonymous):

i thought about it alote :-\ mmm have no clue sry , but since ur talking about Z plane i think if u found one solution , then there is infinitly many solution (have doubt ) ill let u know if i found something else

OpenStudy (anonymous):

updated one year ago!!!

jagr2713 (jagr2713):

After one year, no one has an answer

OpenStudy (anonymous):

i have one constraint... tell me if i am going right re (z) >-1/2

OpenStudy (michele_laino):

please try this substitution: \[\Large z = \rho {e^{i\theta }}\]

OpenStudy (michele_laino):

we can rewrite your equation as follows: \[\Large 1 + z = {z^{\frac{n}{{m + n}}}}\] and after that substitution we have: \[\Large \rho \left( {{e^{i\frac{\pi }{2}}} + {e^{i\theta }}} \right) = {\rho ^{\frac{n}{{m + n}}}}{e^{i\frac{n}{{m + n}}\theta }}\]

OpenStudy (michele_laino):

oops.. \[\Large \rho \left( {{e^{i0}} + {e^{i\theta }}} \right) = {\rho ^{\frac{n}{{m + n}}}}{e^{i\frac{n}{{m + n}}\theta }}\]

OpenStudy (anonymous):

Nice problem.

OpenStudy (ikram002p):

so since |z|<1.... r^n<1 \(\Large \begin{matrix} 1+z=& z^{\frac{n}{n+m}} \\ 1+r(\cos\theta+i \sin\theta ) =& r^{\frac{n}{n+m}}(\cos{\frac{n}{n+m}} \theta + i \sin{\frac{n}{n+m}} )\\ (1+r \cos \theta )+ i(r \sin \theta ) =& (r^{\frac{n}{n+m}} \cos{\frac{n}{n+m}} \theta )+ i (r^{\frac{n}{n+m}}\sin{\frac{n}{n+m}}) \\ \end{matrix} \)

OpenStudy (ikram002p):

\(\begin{matrix} 1+z=& z^{\frac{n}{n+m}} \\ 1+r(\cos\theta+i \sin\theta ) =& r^{\frac{n}{n+m}}(\cos{\frac{n}{n+m}} \theta + i \sin{\frac{n}{n+m}} )\\ (1+r \cos \theta )+ i(r \sin \theta ) =& (r^{\frac{n}{n+m}} \cos{\frac{n}{n+m}} \theta )+ i (r^{\frac{n}{n+m}}\sin{\frac{n}{n+m}}) \\ \end{matrix} \)

OpenStudy (ikram002p):

so as we know imaginary part together , real part together :)

OpenStudy (ikram002p):

\( \Large \begin{matrix} (1+r \cos \theta ) =& (r^{\frac{n}{n+m}} \cos{\frac{n}{n+m}} \theta )\text{ equation 1 } \\ (r \sin \theta ) =& (r^{\frac{n}{n+m}}\sin{\frac{n}{n+m}}\theta )\text{ equation 2 } \\ \end{matrix} \)

OpenStudy (ikram002p):